# Predicting sampling advantage of stochastic Ising Machines for Quantum Simulations

### v1

@@ Abstract
Stochastic Ising machines, sIMs, are highly promising accelerators for optimization and sampling of computational problems that can be formulated as an Ising model. Here we investigate the computational advantage of sIM for simulations of quantum magnets with neural-network quantum states (NQS), in which the quantum many-body wave function is mapped onto an Ising model. We study the sampling performance of sIM for NQS by comparing sampling on a software-emulated sIM with standard Metropolis-Hastings sampling for NQS. We quantify the sampling efficiency by the number of steps required to reach iso-accurate stochastic estimation of the variational energy and show that this is entirely determined by the autocorrelation time of the sampling. This enables predications of sampling advantage without direct deployment on hardware. For the quantum Heisenberg models studied and experimental results on the runtime of sIMs, we project a possible speed-up of 100 to 10000, suggesting great opportunities for studying complex quantum systems at larger scales.
@@ I Introduction
Stochastic Ising Machines (sIM) are an unconventional computational paradigm with high potential to solve certain NP-hard problems, like combinatorial optimization problems, faster and for large problem sizes [1, 2]. To this end, the problem is mapped first onto an Ising model, which is then implemented as physical hardware which solves the Ising model in a massively parallel and energy-efficient way. A particular promising technique to realize sIMs is using probabilistic bits (p-bits), which can fluctuate between 0 and 1 [3, 4]. The advantage of p-bits is that they can be implemented on large scale and operate massively parallel with up to millions of p-bits on a single chip, realizing up to petaflips per second [5].
@@
Given the large range of hardware implementations of sIMs [1], directly benchmarking is challenging even for a single use case and a given NQS architecture. Therefore, in this work we focus on an alternative method to identify potential computational advantage of sIMs for NQS, focusing on the number of steps that a given sampling algorithm needs on both existing digital and reported designs of p-bit computers. We show that the number of steps can be accurately predicted based on the autocorrelation time of the sampling algorithm. We compare both the Metropolis-Hastings (MH) algorithm and Gibbs sampling, which are the workhorse for traditional NQS and sIM implementations, respectively. This approach has the advantage that it provides a hardware agnostic measure of the runtime required that is only defined by the sampling efficiency of the Ising model used to represent quantum states. Evaluation of this approach to pre-trained RBM models and combining the results with reported runtime per step, we find 2 to 4 orders of magnitude speed up compared to single chain MH sampling.
@@ II Model and method
$\psi(s)=\exp\left[\frac{1}{2}\sum_{i=1}^{K}\log{\left(2\cosh{\left(b_{i}+\sum_{j}W_{ij}s_{j}\right)}\right)}\right].$ (3)
where the hidden spins $\{x_{i}\}$ are traced out. For general Hamiltonians $\psi(s)$ needs to be complex, but for the Heisenberg model $\psi(s)$ can be chosen real [25] also allowing real parameters.
We are interested in extracting samples $s$ that are distributed according to $|\psi(s)|^{2}$, directly from an equivalent Ising model. In this case, the Ising model is given by
$H_{\text{ising}}(m)=-\sum_{i=1}^{M+n}h_{i}m_{i}-\sum_{j=1}^{M+n}\sum_{i=1}^{j-1}J_{ij}m_{i}m_{j}$ (4)
and $P(m)=e^{-H_{\text{ising}}(m)}$, where the spins are encoded in a single vector $m=[x_{1}x_{2}\dots x_{M}\ s_{1}\ s_{2}\dots s_{n}]$. The weight matrix $\mathbf{J}$ is given by
$\mathbf{J}=\begin{bmatrix}\mathbf{0}&(\mathbf{W}^{T})_{1\dots M\times 1\dots n}\\ \mathbf{W}_{1\dots n\times 1\dots M}&\mathbf{0}\end{bmatrix}$ (5)
@@
[…]
Figure 1: (a) Restricted Boltzmann Machine with $n$ visible spins and $M$ hidden spins. The visible spins (red) are fully connected to the hidden spins (blue), but there are no intralayer connections. The connections correspond to non-zero weights in the weight matrix $J_{ij}$. (b) Setup of the sampling methods. The MH sampling on the left performs of $N_{\mathrm{MH}}$ MH sweeps each consisting of $n$ antiparallel spin flips of the visible spins. The stochastic Ising Machine (sIM) on the right performs $N_{\mathrm{Ising}}$ sIM sweeps corresponding to updating all hidden and then all visible spins.
@@
$E=\frac{\langle\psi|\hat{H}|\psi\rangle}{\langle\psi|\psi\rangle}=\sum_{s}P(s)E_{\mathrm{loc}}(s)\approx\frac{1}{N}\sum_{m=1}^{N}E_{\mathrm{loc}}(s_{m})$ (7)
where $N$ is the number of samples and $E_{\mathrm{loc}}(s)$ is given by
@@
The states $s_{m}$ are either obtained by MH sampling on the visible spins or sampling the introduced Ising model on an (emulated) sIM, see Fig. 1b.
@@
The sIM samples both hidden and visible spins and uses Gibbs sampling. A key advantage of the RBM’s bipartite structure is that it enables the parallel updating of unconnected spins, namely the visible and hidden (see Fig. 1a). This method, known as chromatic or blocking Gibbs sampling [26, 27], allows for intrinsic parallelism, making it particularly well-suited for deployment on physical hardware.
@@
On the other hand, in the MH sampling traditionally used for NQS, only the visible spins are sampled. Additionally, the MH sampling uses local updates, like a pair of antiparallel spins, which conserves the total magnetization. This is convenient because sampling can be limited to the magnetization 00 sector since the total spin operator $\hat{S}_{\mathrm{tot}}=\sum_{i}\hat{S}^{z}_{i}$ commutes with the Heisenberg Hamiltonian $\hat{H}$, Eq. (1).
@@
Due to these fundamental differences in both sampling space and algorithmic structure, a direct comparison of the two methods is not trivial, and determining which approach is superior depends heavily on the specific context and objectives of the simulation. In the following, the two methods will be compared by looking at the number of steps required to achieve the same accuracy for the variational energy. The accuracy is quantified by the relative error of the variational energy which is defined as
$\varepsilon_{\mathrm{rel}}(N)=\left|\frac{E(N)-E_{\mathrm{b}}}{E_{\mathrm{b}}}\right|$ (9)
@@
where $E(N)$ is the variational energy calculated as a function of number of steps $N$ and $E_{\mathrm{b}}$ serves as a baseline ground state energy. A typical example of such calculations is shown in Fig. 2 for both sIM and MH sampling for the same pre-trained model.
@@
For MH sampling, the number of steps $N$ corresponds to the number of sweeps, where a sweep is defined as attempting to flip pairs of antiparallel spins $n$ times. The baseline energy $E_{\mathrm{b}}$ is obtained using this MH sampling by averaging over $32$ MC chains of each $10.000$ MH sweeps.
For the sIM, the number of steps $N$ corresponds to the number of sIM sweeps, where a sIM sweep is defined as updating all hidden and then all visible spins using chromatic Gibbs sampling. This emulates the hardware realization of a sIM, where bits flip independently with probabilities given by the applied input. Unlike MH sampling, a sIM cannot be constrained to a fixed magnetization sector without distorting the Boltzmann distribution. Therefore, filtering for samples with magnetization 0 afterwards is required but the number of steps $N$ refers to the total amount of steps without filtering.
@@
In order to determine the computational advantage of the sIM, we need to determine the number of steps required for iso-accuracy for the same pre-trained model which is defined as
$\varepsilon_{\mathrm{rel,\text{sIM}}}(N_{\mathrm{sIM}})=\varepsilon_{\mathrm{rel,\text{MH}}}(N_{\mathrm{MH}})\$ (10)
where $N_{\mathrm{sIM}}$ and $N_{\mathrm{MH}}$ represent the number of steps in sIM and MH sweeps, respectively. Additionally, we need the runtime per sIM and MH sweep respectively, which combined with the number of steps, gives the computational advantage expressed as,
$N_{\mathrm{sIM}}t_{\mathrm{S,sIM}}<N_{\mathrm{MH}}t_{\mathrm{S,MH}}$ (11)
where $t_{\mathrm{S,sIM}}$ ($t_{\mathrm{S,MH}}$) refers to the time required per sIM (MH) sweep respectively.
Figure 2: The relative error, $\varepsilon_{\mathrm{rel}}(N)$ as a function of $N$, which corresponds to the number of sIM sweeps or MH sweeps. In (a) and (b) the scaling of the relative error is shown for respectively $n_{\mathrm{spins}}=36$ and $n_{\mathrm{spins}}=484$ spins at $\alpha=2$. The relative error for sIM and MH sampling is calculated as an average value over $10$ runs. The MH fit is a fit of the MH sampling results. The fit is then scaled by the ratio of the autocorrelation time of sIM and MH sampling which should again overlap with the sIM sampling results.
@@
To obtain the number of steps for iso-accuracy, we will now focus on deriving an explicit formula to compare sampling algorithms. In general, the error of an observable $O$ as a function of the number of steps is determined by the autocorrelation time $\tau$ and defined as
$\widetilde{\sigma}_{\overline{O}}=\sqrt{\frac{2\tau}{N\Delta t}\left(\frac{1}{N}\sum_{i=1}^{N}(O_{i}-\overline{O})^{2}\right)}=\sqrt{\frac{2\tau}{N\Delta t}\mathrm{Var}(O)}$ (12)
where $N$ is the number of steps, $\Delta t$ is the interval at which samples are taken and $\mathrm{Var}(O)$ is the variance of observable $O$. The expression is valid when $N\gg 1$ and $\tau\gg\Delta t$ [28, 29].
@@
The relative error $\varepsilon_{\mathrm{rel}}(N)$ scales with the sampling error of the variational energy given by Eq. (12), which decreases when increasing the number of samples until the error becomes on the same order of the baseline. In order to make a practical comparison between the sampling algorithms, we set $\Delta t=1\ \text{MH sweep}$ and $\Delta t=1\ \text{sIM sweep}$ for MH and sIM sampling respectively. Practically, $\Delta t$ can be different to optimize storage and calculate observables, see supplementary C. In the case of $\widetilde{\sigma}_{E}\gg\widetilde{\sigma}_{E_{b}}$, the relative error is then given by
$\varepsilon_{\mathrm{rel}}(N)\sim\widetilde{\sigma}_{\overline{E}}=\sqrt{\frac{2\tau_{\mathrm{}}}{N}\mathrm{Var}(E)}$ (13)
Substituting this result into the iso-accuracy condition, Eq. (10), yields
$\frac{N_{\text{sIM}}}{N_{\text{MH}}}=\frac{\tau_{\text{sIM}}}{\tau_{\text{MH}}}$ (14)
@@
where we used that $\mathrm{Var}(E)$ is the same for both methods. Hence, the number of samples needed for iso-accuracy is determined by the autocorrelation time. Interestingly, for the sIM sampling the number of samples will be related to the sampling efficiency of the Ising model onto which the NQS is mapped.
To verify this analysis, we then fit the relative error of the energy in Fig. 2 to $\varepsilon_{\mathrm{rel}}^{\mathrm{fit}}(N)=a/\sqrt{N}$ for MH sampling (gray dashed line). In addition, we determine the autocorrelation time of both MH and sIM sampling to estimate the scaling of the sIM sampling by
$\varepsilon_{\mathrm{rel,\text{sIM}}}(N)=\sqrt{\frac{\tau_{\text{sIM}}}{\tau_{\text{MH}}}}\ \varepsilon_{\mathrm{rel,\text{MH}}}^{\mathrm{fit}}(N).$ (15)
which is indicated by the black dashed line. Indeed, it is found that the dashed line fits the slope of the sIM sampling well.
@@
In the following, we directly use Eq. (14) to estimate the number of steps required on a sIM as compared to MH sampling by measuring the autocorrelation of sIM and MH sampling. To this end, we consider pre-trained RBM models which are optimised for the ground state of the 2D ($L\times L$) antiferromagnetic Heisenberg model with stochastic reconfiguration [9, 30], see supplementary B.1 and B.2. Each pre-trained RBM model has a corresponding Ising model, with a given exchange matrix and bias vector, Eq. (5) and (6) respectively. We constructed a data set of several these pre-trained models with various combinations of $\alpha$ and $n_{\mathrm{spins}}$ ($=n=L^{2}$). This enables a systematic analysis of the scaling advantages of sIM sampling as function of the representational power and system size. For each combination of $\alpha$ and $n_{\mathrm{spins}}$, $5$ models were trained and for each model the sIM and MH autocorrelation time was measured.
@@
Figure 3: Ratio of autocorrelation time $\tau_{\mathrm{sIM}}/\tau_{MH}=N_{\mathrm{sIM}}/N_{MH}$ for different models. Each point is the average of $5$ pre-trained models with the same $\alpha$ and $n_{\mathrm{spins}}=L^{2}$.
@@
The results at $\alpha=1$ were excluded from the figure because they are less accurate [9, 10], moreover extensive hyperparameter tuning is required to consistently find an accurate ground state for this value. The autocorrelation time per model was measured for $5$ ($10$) chains for sIM (MH) sampling, see supplementary C.
@@
The results for the predicted ratio $N_{\mathrm{sIM}}/N_{\mathrm{MH}}$ are shown in Fig 3. It is clear that the ratio of the autocorrelation time of sIM and MH can differ orders of magnitude. Interestingly, models with $\alpha=2$ have a significantly lower autocorrelation time compared to the models with $\alpha>3$, which we elaborate on in the discussion below.
@@
Figure 4: Performance comparison for generating the same amount of independent samples as one MH-sweep across five methods as a function of $n_{\mathrm{spins}}=L^{2}$ for $\alpha=2$. The baseline, MH (CPU) obtained using UltraFast [31], is a traditional NQS implementation with MH sampling. Furthermore, there are four sIM implementations: CPU-based implementation, FPGA running at 70 MHz [32], conservative projection based on [33], and optimistic projection based on [5].
@@
Equipped with the ratio of the autocorrelation time, we focus on projecting the possible computational advantage of deployment on sIMs. To this end, we focus on the $\alpha=2$ case and consider reported values of the runtime per step on physical sIMs. The projections represent the time required to generate the same amount of independent samples as one MH sweep, which is calculated from the ratio of the number of steps, Eq. (14), and the runtime per step on the hardware. On all (hardware) implementations, a single MC chain is assumed. In Fig. 4, the MH (CPU) baseline, is the performance of UltraFast [31] on a machine with dual AMD EPYC 7601 CPU. The sIM (CPU) corresponds to a chromatic Gibbs sampling code written in Julia on the same machine. This implementation uses multithreading for matrix-vector multiplication. Benchmarks for a (single-chain) sIM GPU implementation on an NVIDIA L40S GPU are omitted, as it only outperforms the sIM CPU implementation for systems with $n_{\mathrm{spins}}>700$ spins, which is larger than the system sizes considered here, see supplementary D.
@@
The conservative sIM projection is based on a device combining the IBM HERMES Project Chip with stochastic MTJs [33]. This chip design features 64 analog crossbars of size $256\times 256$ for MVMs with a latency of 133 ns [33]. The typical latency of the stochastic MTJs is $0.1-10$ ns [34, 5]. Therefore, we assume 400 ns per IM sweep for this device, also incorporating time required for splitting the MVM over multiple cores.
The optimistic sIM projection follows the analysis of [5] for a projected asynchronous sIM which is based on a $100\times 100$ crossbar with a latency of $10$ ps [35]. Again combining this with the stochastic MTJs, we assume a latency of $4$ ns per sIM sweep, or 2 orders of magnitude faster than the conservative projection.
From Fig. 4, it is clear that the CPU based sIM implementation performs similar to the MH (CPU, UltraFast) implementation. The conservative projection is 2 orders of magnitude faster than the baseline and demonstrates a more favorable scaling with system size. Interestingly, the FPGA implementation is faster than conservative projection. Lastly, the optimistic projection is 2 orders of magnitude faster than the conservative projection and slightly faster than the FPGA implementation.

### v2

@@ Abstract
Stochastic Ising machines, sIMs, are highly promising accelerators for optimization and sampling of computational problems that can be formulated as an Ising model. Here we investigate the computational advantage of sIM for simulations of quantum magnets with neural-network quantum states (NQS), in which the quantum many-body wave function is mapped onto an Ising model. We study the sampling performance of sIM for NQS by comparing sampling on a software-emulated sIM with standard Metropolis-Hastings sampling for NQS. We quantify the sampling efficiency by the number of computational steps required to reach iso-accurate stochastic estimation of the variational energy and show that this is entirely determined by the autocorrelation time of the sampling. This enables predictions of sampling advantage without direct deployment on hardware. Although sampling of the quantum Heisenberg models studied exhibits much longer autocorrelation times on sIMs, the massively parallel sampling of hardware sIMs leads to a projected speed-up of 100 to 10000, suggesting great opportunities for studying complex quantum systems at larger scales.
@@ I Introduction
Stochastic Ising Machines (sIM) are an unconventional computational paradigm with high potential to solve certain NP-hard problems, like combinatorial optimization problems, faster and for large problem sizes [1, 2]. To this end, the problem is mapped first onto an Ising model, which is then implemented as physical hardware which solves the Ising model in a massively parallel and energy-efficient way. A particularly promising technique to realize sIMs is using probabilistic bits (p-bits), which can fluctuate between 0 and 1 [3, 4]. The advantage of p-bits is that they can be implemented on large scale and operate massively parallel with up to millions of p-bits on a single chip, realizing up to petaflips per second [5].
@@
Given the large range of hardware implementations of sIMs [1], directly benchmarking is challenging even for a single use case and a given NQS architecture. Therefore, in this work we focus on an alternative method to identify potential computational advantage of sIMs for NQS, focusing on the number of computational steps that a given sampling algorithm needs on both existing digital and reported designs of p-bit computers. We show that the number of steps can be accurately predicted based on the autocorrelation time of the sampling algorithm. We compare both the Metropolis-Hastings (MH) algorithm and Gibbs sampling, which are the workhorse for traditional NQS and sIM implementations, respectively. This approach has the advantage that it provides a hardware agnostic measure of the runtime required that is only defined by the sampling efficiency of the Ising model used to represent quantum states. Evaluation of this approach to pre-trained RBM models and combining the results with reported runtime per step, we find 2 to 4 orders of magnitude speed up compared to single chain MH sampling.
@@ II Model and method
$\psi(s)=\exp\left[\frac{1}{2}\sum_{{j}=1}^{{m}}\log{\left(2\cosh{\left(b_{{j}}+\sum_{{i=1}}^{{n}}W_{ij}s_{{i}}\right)}\right)}\right].$ (3)
where the hidden spins $\{x_{{j}}\}$ are traced out. For general Hamiltonians $\psi(s)$ needs to be complex, but for the Heisenberg model $\psi(s)$ can be chosen real [24] also allowing real parameters.
We are interested in extracting samples $s$ that are distributed according to $|\psi(s)|^{2}$, using an equivalent Ising Model of which the Hamiltonian is given by
$H_{\text{ising}}({z})=-\sum_{i=1}^{{m}+n}h_{i}{z}_{i}-\sum_{j=1}^{{m}+n}\sum_{i=1}^{j-1}J_{ij}{z}_{i}{z}_{j}$ (4)
and $P(z)=e^{-H_{\text{ising}}(z)}$}, where the spins are encoded in a single vector ${z}=[x_{1}x_{2}\dots x_{{m}}\ s_{1}\ s_{2}\dots s_{n}]$. Note that the connectivity of the Ising Hamiltonian here is all-to-all, which differs from the canonical nearest-neighbor Ising model typically used in physics. The weight matrix $\mathbf{J}$ is given by
$\mathbf{J}=\begin{bmatrix}\mathbf{0}&(\mathbf{W}^{T})_{{m\times n}}\\ \mathbf{W}_{{n\times m}}&\mathbf{0}\end{bmatrix}$ (5)
@@
$\mathbf{h}=[\mathbf{b}_{{m}}\ \mathbf{0}_{{n}}].$ (6)
Figure 1: (a) Restricted Boltzmann Machine with $n$ visible spins and m{m} hidden spins. The visible spins (red) are fully connected to the hidden spins (blue), but there are no intralayer connections. The connections correspond to non-zero weights in the weight matrix $J_{ij}$. (b) Setup of the sampling methods. The MH sampling on the left performs of $N_{\mathrm{MH}}$ MH sweeps each consisting of $n$ antiparallel spin flips of the visible spins. The stochastic Ising Machine (sIM) on the right performs $N_{\mathrm{sIM}}$ sIM sweeps which corresponds to the number of samples in the magnetization zero sector.
@@
$E=\frac{\langle\psi|\hat{H}|\psi\rangle}{\langle\psi|\psi\rangle}=\sum_{s}P(s)E_{\mathrm{loc}}(s)\approx\frac{1}{N}\sum_{{k}=1}^{{M}}E_{\mathrm{loc}}(s_{{k}})$ (7)
where $M$ is the number of samples and $E_{\mathrm{loc}}(s)$ is given by
@@
The states $s_{{k}}$ are either obtained by MH sampling on the visible spins or sampling the introduced Ising model on an (emulated) sIM, see Fig. 1b.
@@
The sIM samples both hidden and visible spins and uses Gibbs sampling. A key advantage of the RBM’s bipartite structure is that it enables the parallel updating of unconnected spins, namely the visible and hidden, see Fig. 1a. This method, known as chromatic or blocking Gibbs sampling [25, 26], allows for intrinsic parallelism, making it particularly well-suited for deployment on physical hardware.
@@
On the other hand, in the MH sampling traditionally used for NQS, only the visible spins are sampled. Additionally, the MH sampling uses local updates, like a pair of antiparallel spins, which conserves the total magnetization, $\sum_{i=1}^{n}s_{i}$. This is convenient because sampling can be limited to the magnetization zero sector since the total spin operator $\hat{S}_{\mathrm{tot}}=\sum_{i=1}^{n}\hat{S}^{z}_{i}$ commutes with the Heisenberg Hamiltonian $\hat{H}$, Eq. (1).
@@
Due to these fundamental differences in both sampling space and algorithmic structure, a direct comparison of the two methods is not trivial, and determining which approach is superior depends heavily on the specific context and objectives of the simulation. In the following, the two methods will be compared by looking at the number of Monte Carlo steps $N$ required to achieve the same accuracy for the variational energy. The accuracy is quantified by the relative error of the variational energy which is defined as
$\varepsilon(N)=\left|\frac{E(N)-E_{\mathrm{b}}}{E_{\mathrm{b}}}\right|$ (9)
@@
where $E(N)$ is the variational energy calculated and $E_{\mathrm{b}}$ serves as a baseline ground state energy.
@@
The latter is obtained using MH sampling by averaging over $32$ MC chains of $100.000$ MH sweeps, which is defined as attempting to flip pairs of antiparallel spins (MH flip) $n$ times, see Fig. 1b.
@@
In order to determine the computational advantage of the sIM, we need to determine the number of steps $N$ required for iso-accuracy for the same pre-trained model which is defined as
$\varepsilon_{\mathrm{\text{sIM}}}(N_{\mathrm{sIM}})=\varepsilon_{\mathrm{\text{MH}}}(N_{\mathrm{MH}})\\$ (10)
where the subscripts sIM and MH indicate the values of $N$ and $\varepsilon$ for sIM and MH sampling, respectively.
The computational advantage is identified when the total runtime of sIM sampling is less than that of MH sampling,
$N_{\mathrm{sIM}}T_{\mathrm{sIM}}<N_{\mathrm{MH}}T_{\mathrm{MH}}.$ (11)
where $T_{\mathrm{sIM}}(T_{\mathrm{MH}})$ denotes the runtime per step of sIM (MH) step.
@@
To obtain the number of steps for iso-accuracy, we will now focus on deriving an explicit formula to compare sampling algorithms. Since both the sIM and MH are Markov chain Monte Carlo (MCMC) methods, they generate correlated samples. For this case, the sampling error for an observable $O$ is given by
$\sigma_{\overline{O}}(N,\rho)=\sqrt{\frac{1+2\sum_{l=1}^{N}\rho(l)}{N}\mathrm{Var}(O)}$ (12)
in the limit of large $N$, where $\rho(l)$ is the autocorrelation function and $\mathrm{Var}(O)$ is the variance of observable $O$. The autocorrelation function decays exponentially with the autocorrelation time $\tau$: for large $l$, $\rho(l)\sim e^{-l/\tau}$ [27, 28, 29].
This behavior is shown in Fig. 2a for both MH and sIM sampling as a function of the number of sweeps (for sIM and MH) and also in flips (for MH). Here a sIM sweep is defined as performing sIM updates until a new configuration with zero net magnetization is obtained. A sIM update consists of updating all hidden and then all visible spins using chromatic Gibbs sampling. This resembles the hardware realization of a sIM, where bits flip independently with probabilities given by the applied input. Unlike MH sampling, a sIM cannot be constrained to a fixed magnetization sector without distorting the desired Boltzmann distribution. Since observables are only defined for samples with zero magnetization, we discard samples with non-zero magnetization.
Due to the correlated nature of MCMC, it is more efficient in practice to evaluate observables at a sampling interval $\Delta t=N/M>1$, where $M$ is the number of samples used in the calculations. For large $M$ and by scaling $\tau$ in units of $\Delta t$, we can approximate $\sum_{k=1}^{M}\rho(k)\approx\sum_{k=1}^{M}e^{-k/(\Delta t/\tau)}\approx 1/(e^{\Delta t/\tau}-1)$.
@@
The relative error $\varepsilon({M})$ scales with the sampling error of the variational energy given by Eq. (12), which decreases when increasing the number of samples until the error becomes on the same order of the baseline. In the case of $\sigma_{E}\gg\sigma_{E_{b}}$ and large $M$, the relative error is then given by
$\varepsilon(M)\sim\sigma_{\overline{E}}(M,\Delta t/\tau)=\sqrt{\frac{1+2/(e^{\Delta t/\tau}-1)}{M}\mathrm{Var}(E)}.$ (13)
Figure 2: (a) Autocorrelation function as a function of $l$ for $n=36$, $\alpha=2$. Bottom axis shows the interval in units of the sweep size. The top axis gives the number of MH flips, which only applies to MH sampling. The measured autocorrelation time is $\tau_{\mathrm{MH}}=0.32\pm 0.02$ MH sweeps and $\tau_{\mathrm{sIM}}=2.01\pm 0.01$ sIM sweeps. On the top x-axis, the MH sweeps is converted into MH flips. (b) The relative error, $\varepsilon(M)$ as a function of $M$ for $n=36,64,100,144,400,484$ spins at $\alpha=2$. Here $M$ corresponds to the number of samples at a sampling interval of $\Delta t_{\mathrm{sIM,MH}}=\tau_{\mathrm{sIM,MH}}$ respectively. The relative error for sIM and MH sampling is calculated as a mean value over $32$ runs and the error of the relative error indicate the standard error of the mean. The fit highlights the scaling with $M^{-1/2}$. The shaded error bars signify one standard deviation estimated from $32$ chains.
In Fig. 2b, sampling with $\Delta t=\tau$ for sIM and MH gives the same scaling for the relative error which is confirmed by the fitted black dashed line $\varepsilon_{\mathrm{fit}}(M)=aM^{-1/2}$. This confirms the scaling in Eq. (13) and that shows sampling at the fixed interval $\Delta t=\tau$ and given number of samples $M$ yields the same relative error for both methods.
To estimate the computational cost, we want to know the number of steps $N$ required to reach iso-accuracy. For that we want to take the same number of samples from both samplers and we set $M_{\mathrm{sIM}}=M_{\mathrm{MH}}$ and substitute this into the iso-accuracy condition, Eq. (10), which yields
$\frac{\tau_{\mathrm{sIM}}}{\tau_{\mathrm{MH}}}=\frac{\Delta t_{\mathrm{sIM}}}{\Delta t_{\mathrm{MH}}}$ (14)
@@
where we used that $\mathrm{Var}(E)$ is the same for both methods. Hence, the sampling interval and therefore the number of steps needed for iso-accuracy is determined by the autocorrelation time. Interestingly, for the sIM sampling the number of steps will be related to the sampling efficiency of the Ising model onto which the NQS is mapped.
@@
In the following, we directly use Eq. (14) to estimate the number of steps required on a sIM as compared to MH sampling by measuring the autocorrelation of sIM and MH sampling. To this end, we consider pre-trained RBM models which are optimised for the ground state of the 2D ($L\times L$) antiferromagnetic Heisenberg model with stochastic reconfiguration [8, 30], see Appendix B.1. We also compared the accuracy to QMC [31] in Appendix B.2. Each pre-trained RBM model has a corresponding Ising model, with a given exchange matrix and bias vector, Eq. (5) and (6) respectively. We constructed a data set of several these pre-trained models with various combinations of $\alpha$ and n{n} ($L^{2}$). This enables a systematic analysis of the scaling advantages of sIM sampling as function of the representational power and system size. For each combination of $\alpha$ and $n$, 55 models were trained and for each model the sIM and MH autocorrelation time was measured.
@@
Figure 3: Ratio of autocorrelation time $\tau_{\mathrm{sIM}}/\tau_{MH}=\Delta t_{\mathrm{sIM}}/\Delta t_{MH}$ for different models. Each point is the average of $5$ pre-trained models with the same $\alpha$ and $n=L^{2}$.
@@
The results at $\alpha=1$ were excluded from the figure because they are less accurate [8, 9], moreover extensive hyperparameter tuning is required to consistently find an accurate ground state for this value. The autocorrelation time per model was measured for 55 ($10$) chains for sIM (MH) sampling, see Appendix C.
@@
The results for the ratio $\tau_{\mathrm{sIM}}/\tau_{\mathrm{MH}}$ are shown in Fig 3. It is clear that the ratio of the autocorrelation time of sIM and MH can differ by orders of magnitude. Interestingly, models with $\alpha=2$ have a significantly lower autocorrelation time compared to the models with $\alpha>3$, which we elaborate on in the discussion below.
@@
Figure 4: Performance comparison for generating one sample across five methods as a function of ${n}=L^{2}$ for $\alpha=2$. The baseline, MH (CPU) obtained using UltraFast [32], is a traditional NQS implementation with MH sampling. Furthermore, there are four sIM implementations: CPU-based implementation, FPGA running at 70 MHz [33], conservative projection based on [34], and optimistic projection based on [5].
Equipped with the ratio of the autocorrelation time, we focus on projecting the possible computational advantage of deployment on sIMs. To this end, we focus on the $\alpha=2$ case and consider reported values of the runtime per step on physical sIMs. Since NQS literature often reports the number of required samples at a sampling interval of one MH sweep, we will report the time to generate the same effective sample for MH and sIM. The time per sample on the sIM is calculated from the ratio of the sampling intervals, Eq. (14), with $\Delta t_{\mathrm{MH}}=1\mathrm{MH\ sweep}$ and the runtime per update on the sIM hardware:
$\text{Time per sample}=\Delta t_{\mathrm{sIM}}T_{\mathrm{sIM}}=\frac{\tau_{\mathrm{sIM}}}{\tau_{\mathrm{MH}}f}T_{\mathrm{sIM}}^{\mathrm{update}}$ (15)
where $\tau_{\mathrm{sIM}}$, $\tau_{\mathrm{MH}}$, $f$ are all dependent on the number of spins $n$ and $\alpha$. Since the sIM is not constrained to the magnetization zero sector, the fraction $f$ enters, which is the ratio of the number of samples with zero magnetization to the total number of samples: $T_{\mathrm{sIM}}=T_{\mathrm{sIM}}^{\mathrm{update}}/f$. The MH sampling baseline on CPU, $\text{Time per sample}=T_{\mathrm{MH}}$, is directly measured in our experiments and is also dependent on $n$ and $\alpha$. On all (hardware) implementations, a single MC chain is assumed.
@@
In Fig. 4, the MH (CPU) baseline, is the performance of UltraFast [32] on a machine with dual AMD EPYC 7601 CPU. The sIM (CPU) corresponds to a chromatic Gibbs sampling code written in Julia on the same machine. This implementation uses multithreading for matrix-vector multiplication [35]. Benchmarks for a (single-chain) sIM GPU implementation on an NVIDIA L40S GPU are omitted, as it only outperforms the sIM CPU implementation for systems with n>700{n}>700 spins, which is larger than the system sizes considered here, see Appendix E.
@@
The conservative sIM projection is based on a device combining the IBM HERMES Project Chip with stochastic MTJs [34]. This chip design features 64 analog crossbars of size $256\times 256$ for MVMs with a latency of 133 ns [34]. We assume that the typical latency of the stochastic MTJs is $2$ ns [37, 5]. Implementing chromatic Gibbs sampling requires two MVMs, splitting the MVM over multiple cores (if MVM is bigger than one crossbar) and two read-outs of the MTJs. Therefore, we estimate 400 ns per sIM sweep for this device, see Appendix F for more details.
The optimistic sIM projection follows the analysis of [5] for a projected asynchronous sIM which is based on a $100\times 100$ crossbar with a latency of $10$ ps [38]. Again combining this with the stochastic MTJs, we estimate a latency of $4$ ns per sIM sweep, or 2 orders of magnitude faster than the conservative projection, see Appendix F.
From Fig. 4, it is clear that the CPU based sIM implementation performs similar to the MH (CPU, UltraFast [32]) implementation. The conservative projection is 2 orders of magnitude faster than the baseline and demonstrates a more favorable scaling with system size. Interestingly, the FPGA implementation is faster than conservative projection. Lastly, the optimistic projection is 2 orders of magnitude faster than the conservative projection and slightly faster than the FPGA implementation.